A 7-foot hole is dug that measures 26 feet by 26 feet. If a truck with a capcity of 6 cubic yards is used to haul away the dirt, how many trips must the truck make?
Found 2 solutions by checkley71, edjones:
Answer by checkley71(8403)

7*26*26=4732 cubic feet of dirt.
6 cubic yards=6*27=162 cubic feet which one truck load.
4732/162=29.2 truck loads will be required to haul away all the dirt.

Answer by edjones(8007)

(26/3)^2*7/3=175.259 yds^3 divide ft by 3 to get yards.
175.259/6=29.2099 truckloads
=30 trips
.
Ed